# Data Release 1 of the Dark Energy Spectroscopic Instrument

### v1

@@ F.2 Milky Way Survey VACs
MWS VAC provides analysis of stellar spectra by the Milky Way Survey Working Group (Koposov et al. 2025, in preparation). The analysis involves fitting DESI spectra using two pipelines: RVSpecFit (Koposov et al., 2024) and FERRE (Allende-Prieto & Apogee Team, 2023) (see Cooper et al. 2023 and Koposov et al. 2024 for more details). The RVSpecFit pipeline derives radial velocity measurements, stellar parameters, and abundances, while FERRE provides stellar parameters and abundance measurements. Both pipelines have been applied to approximately 6.5M coadded spectra of stars or possible stars from DESI DR1. Additionally, the RVSpecFit pipeline was run on approximately 9.5M single-exposure spectra, enabling radial velocity variability analyses. Since the DESI EDR, both pipelines have been significantly improved, with RVSpecFit now using neural networks to interpolate stellar spectra, thus eliminating issues with clustering of stellar parameters on grid nodes that was present in the EDR VAC. The data products in this VAC include tables of stellar parameters, radial velocities, their uncertainties, and the best-fit models for the DESI spectra. In addition, the data tables have been cross-matched with the latest Gaia DR3 release.
@@
MWS BHB VAC contains spectroscopically confirmed BHB stars. BHB stars are excellent tracers of the Milky Way halo; they are intrinsically bright, so they can be observed to large distances, and there exists a simple polynomial relationship between their color and absolute magnitude (Deason et al., 2011; Belokurov & Koposov, 2016), meaning that their distances can be computed with high precision. BHB candidates were targeted in the MWS (see §4.4.4 in Cooper et al., 2023) and as secondary targets observed in dark time (see Myers et al., 2023a). Due to photometric uncertainties, targeted BHB candidates contain contamination primarily from more nearby blue straggler stars. The DESI spectroscopic observations, however, allow us to remove this contamination. The BHB VAC contains over 6,300 spectroscopically confirmed BHB stars, all having distances derived from photometry, as described in Byström et al. (2024). This BHB catalogue contains the stellar parameters from the RVSpecFit pipeline (Koposov et al., 2024), including radial velocity, surface gravity and effective temperature, and reaches heliocentric distances of about 120 kpc.
@@
MWS SpecDis VAC provides distance measurements for over 4M stars from DR1 following the selection criteria outlined in Li et al. (2025). The VAC is based on a data-driven approach which adopts a neural network to establish the connection between stellar spectra and luminosity. It is trained on Gaia parallax and Gaia -band apparent magnitude measurements, with the following key improvements: (1) We adopt a special training label to avoid cutting off negative parallaxes; (2) We do not include any parallax uncertainty cuts for the training sample, but we include the parallax uncertainties in the loss function. This approach ensures that the training sample does not suffer from biases in parallax, and our training sample includes a significant number of distant giants. We can achieve significant improvements in our measurements than Gaia parallaxes beyond 7 kpc. (3) We adopt principal component analysis to decrease the noise and the dimensionality of the spectra. In general, the distance uncertainties decrease with increasing average signal-to-noise ratio in the B and R arms of the DESI stellar spectra (see Table 2). We adopt a Gaussian mixture model to identify candidate binary stars. The final catalog offers distance, distance uncertainty, the binary candidate flag plus other useful photometric and astrometric information, cross-matched from the RVSpecFit values reported in the MWS VAC (Koposov et al. 2025, in preparation) and Gaia DR3.
@@
SPDist VAC provides spectrophotometric distances in the form of absolute magnitude in the Gaia -band for all stars observed by the DESI MWS, using a data-driven approach similar to Thomas & Battaglia (2022). Distances are estimated using a fully connected multi-layer perceptron (MLP) model using Gaia DR3 photometry (Gaia Collaboration et al., 2023), spectroscopic metallicity from the SP pipeline, and atmospheric parameters (effective temperature and surface gravity) derived from a combination of values from both the RV and SP pipelines (Koposov et al., 2024) as inputs. The MLP outputs the absolute magnitude in the Gaia -band, which is subsequently used to calculate distances. Some stars exhibit significant discrepancies in spectroscopic parameters derived by the RV and SP pipelines, likely leading to unreliable distance estimates. These suspicious stars are flagged with the FLAG_GOOD parameter. We recommend using this flag along with the quality flags recommended for the RV and SP pipelines (Koposov et al., 2024). For distance modulus computations, we recommend using the median of the absolute magnitude distribution (MG_50), while uncertainties can be derived from the 16th and 84th percentiles of the distribution. These uncertainties should be combined in quadrature with the intrinsic precision of the method (0.167 mag, 8% relative distance precision) to obtain the total uncertainty. The primary difference between the methodologies of SpecDist and SPdist is that SpecDist predicts distances from the stellar spectra, whereas SPdist predicts distances from a full list of stellar parameters by the MWS RV and SP pipelines. A forthcoming paper will present a detailed comparison of the two methods.
@@
Stellar Reddening VAC contains the observed spectra and the RVSpecFit model spectra of stars that provided reddening measurements for the DESI dust map (Zhou et al., 2024). The DESI dust map, publicly released with Zhou et al. (2024), is used in the DESI DR1 cosmology analysis (DESI Collaboration et al., 2024c). Also included in this VAC is the stellar catalog containing the per-star reddening measurements in the Dark Energy Camera (DECam; Flaugher et al., 2015) filters as well as other properties from the DESI pipeline and the imaging catalogs. While the paper uses the first two years of DESI data, only the DR1 portion of the data is released in the VAC. Also note that while both the MWS VAC and this VAC use RVSpecFit and are based on the same DESI data, there are some key differences: (1) the RVSpecFit model spectra in this VAC are the zero-extinction spectra, whereas the MWS VAC model spectra include an additional multiplicative term to fit the smooth-varying component of the observed spectra; and (2) the DR1 MWS VAC uses a different (newer) version of RVSpecFit than the version used here. For general purposes (other than for studying dust), the MWS VAC should be used instead of this VAC.
@@ F.3 Extragalactic Science VACs
HETDEX VAC contains HETDEX (Hill et al., 2021) and DESI spectra of HETDEX-selected Ly emitter candidates followed up by DESI. The VAC also contains the emission line fits from both sets of spectra. Details of the analysis can be found in Landriau et al. (2025).
@@
DESIVAST VAC contains cosmic voids identified within the DESI DR1 volume. The void locations are computed using a volume-limited subsample of the BGS Bright sample (see Hahn et al. 2023) extending to . We apply evolutionary corrections to 479,486 galaxies and enforce a magnitude cut of . We then apply three void-finding algorithms: VoidFinder (El-Ad & Piran, 1997; Hoyle & Vogeley, 2002; Douglass et al., 2022), V2/VIDE (Neyrinck, 2008; Sutter et al., 2015), and V2/REVOLVER (Nadathur et al., 2019), to obtain three comparable void catalogs. Our VoidFinder catalog contains 3,765 voids, with 1,489 non-edge voids defined as those not bordering on the survey edges. For V2/VIDE (V2/REVOLVER), we find 1,478 (1,992) voids with 297 (389) non-edge voids. Further information regarding this VAC can be found in Rincon et al. (2024).
@@
Dwarf Galaxy VAC presents a sample of extragalactic dwarf galaxies () identified in DESI DR1. The catalog includes galaxies selected from the BGS and ELG samples, as well as the LOW-Z secondary target program (Darragh-Ford et al., 2023), spanning a redshift range of . However, due to differences in target selection criteria across these samples, the catalog is not uniformly selected. Stellar masses are derived using CIGALE494949https://cigale.lam.fr (Boquien et al., 2019; Siudek et al., 2024) and optical color-based prescriptions (Mao et al., 2024; de los Reyes et al., 2024). In a forthcoming paper (Manwadkar et al. 2025, in preparation), we will supplement this catalog with information on spectroscopic and photometric completeness, and an analysis of quenched fractions as a function of environment.
@@
EmFit VAC provides emission-line fitting results using the EmFit code (Pucha et al., 2024) run on low-redshift () galaxy spectra, focusing on eight nebular emission lines and doublets: H, [O III] 4959,5007, [N II] 6548,83, H, and [S II] 6716,31. The code tests for the presence of extra components in the [S II] and [O III] doublets independently, and the profiles of the rest of the narrow components are designed to match the profile of the [S II] lines. It further tests for the presence of a possible broad component in the Balmer lines. The flux and width measurements along with their uncertainties for all detected components are reported in this catalog (Pucha et al. 2025, in preparation).
@@
FastSpecfit VAC delivers a broad range of observed-frame, rest-frame, and intrinsic physical parameters for all extragalactic () targets in DR1 observed by DESI (Moustakas 2023; Moustakas et al. 2025, in preparation).505050https://fastspecfit.readthedocs.io It uses physically motivated stellar population synthesis and emission-line templates to model each DESI spectrum jointly with the optical through infrared broadband photometry. The measurements included in this VAC include: stellar velocity dispersions; stellar masses; K-corrections and rest-frame colors and magnitudes; emission-line velocity widths, velocity shifts, fluxes, and equivalent widths (separately both narrow and broad emission-line components); and much more.
@@
Extended Halo-based Group VAC is the updated version of the extended halo-based group catalog derived from the Legacy Surveys DR9 for a galaxy sample limited to -band apparent magnitude . For all the DESI target classes included in DR1, the photometric redshifts will be updated to spectroscopic redshifts. Yang et al. (2021) provide more details regarding the extended halo-based group finder and sample selection.
@@
Mass EMLines VAC provides stellar mass and emission line measurements for all galaxies in DESI DR1 with reliable redshift measurements. Stellar masses are derived using CIGALE (Boquien et al., 2019), which employs the broadband -, -, -, -, and -band photometry from the Legacy Surveys, and spectrophotometry of 10 artificial bands generated through convolution with DESI spectra. A main set of optical emission lines are measured by a single Gaussian fit, with absorption correction through stellar continuum fitting performed using STARLIGHT515151http://www.starlight.ufsc.br (Cid Fernandes et al., 2005, 2011). Additionally, the catalog includes stellar population properties derived by CIGALE and those obtained by STARLIGHT using DESI spectra. For additional information see Zou et al. (2024) and Zou et al. (2025, in preparation).
@@
Strong Lensing VAC provides a comprehensive catalog of spectroscopic observation of strong gravitational lenses observed by DESI. The observations are primarily obtained through a program that seeks to obtain spectroscopic redshifts for both the lensing galaxies and the lensed sources for 1800 new strong lensing candidate systems (Huang et al. 2025, in preparation; Storfer et al. 2025, in preparation). Objects included in the catalog were visually inspected in order to make a determination on the quality of the Redrock redshift, and to determine the nature of the system. Systems with at least one confirmed redshift for either the lens or source are included in the VAC. In addition to relevant data from the default DESI pipeline, the VAC includes data from the FastSpecFit VAC, including velocity dispersion, star-formation rate, and stellar mass. Also included is diameter aperture photometry at the location of the fibers targeting lenses and sources. Lastly, with a number of Ly emitters and absorbers not currently fit for redshifts via Redrock, we use custom template-fitting within the Redrock framework to fit for our own redshifts. Systems that are assigned redshifts this way are noted in the catalog.
@@ F.4 Quasar Science VACs
AGN/QSO Summary VAC provides AGN and QSO identification for all target classes observed within DR1 for objects with at least one DESI spectrum. Identification of QSOs and AGN is achieved through multiple means, including optical and ultraviolet emission-line diagnostics and multi-wavelength photometry. The VAC enables rapid selection of more complete versus more pure methodologies, as well as selection through any one AGN diagnostic (Juneau et al. 2025, in preparation). A tutorial accompanies the catalog which demonstrates both how to use the catalog for rapid AGN / QSO selection as well as using the open source diagnostic code to select a specific line diagnostic and to choose the dividing line methodologies of those diagnostics. An example might be to test the [S II] Baldwin et al. (1981) classification assuming the more recent dividing lines of Law et al. (2021, implemented by default) compared with the traditional dividing lines of Kewley et al. (2001), which were reported to potentially overestimate the presence of AGN (e.g., Pucha et al., 2024).
@@
AGN Host Properties VAC provides physical properties for approximately 17M galaxies derived through spectral energy distribution fitting using CIGALE (v.22.1; Boquien et al., 2019). This analysis accounts for contributions from both stars and, if present, an AGN. CIGALE uses the principle of energy balance, where dust-absorbed stellar emission in the ultraviolet and optical bands is re-emitted in the infrared; it estimates key galaxy and AGN properties such as stellar mass, star formation rate, and the AGN fraction (the contribution of the AGN dusty torus to the total IR luminosity (e.g., Ciesla et al., 2015; Salim et al., 2018; Yang et al., 2020, 2022). A detailed description of the catalog and its statistical properties based on data in the DESI EDR can be found in Siudek et al. (2024), with additional follow-up analysis using spectral energy distribution modeling to identify AGN to be described in Siudek et al. (2025, in preparation).
@@
BHMass VAC provides Mg II-based supermassive black hole mass estimates for 490,648 quasars at from DESI DR1. Approximately 35% of the sample have mass uncertainties less than 0.5 dex. The iron-corrected mass is estimated using the fitting process and estimator introduced by Pan et al. (2025), which accounts for the effects of the Eddington ratio. Additionally, we integrate fitting results from the FastSpecFit VAC and offer alternative mass estimates based on the formulations by Shen et al. (2011); Le et al. (2020); Yu et al. (2023), providing users with flexibility in choosing their preferred estimator. This catalog will be periodically updated and extended to include more DESI quasars.
@@
DLA NN and GP Finder VAC contains the catalog of DLAs used for the DR1 Ly BAO analysis. The catalog results from the combination of two DLA finding algorithms. The first one is based on a convolution neural network trained with mock DESI spectra whose algorithm and performance is described in Wang et al. (2022). The second algorithm, which uses a Gaussian Process (GP), is based on the GP model of Ho et al. (2020). Both outputs from these algorithms are retained in the combined catalog and DLA candidates are merged when they are found by both methods with a velocity difference of less than 800 km s-1. The merged catalog includes estimated redshifts and H I column densities together with confidence flags from both algorithms. It contains 54,416 candidate DLAs with H I column density cm-2 within Ly forest spectra with signal-to-noise ratio larger than three.
@@
DLA Template Finder VAC contains candidate DLAs identified with the DLA Toolkit software, which uses spectral template-fitting. The catalog was constructed from quasar spectra with , allowing for a maximum of three detections per sight-line; it contains an estimated DLA redshift, HI column density, and detection significance for each DLA. Brodzeller et al. (2025, in preparation) provide the full details on the catalog’s construction, parameter accuracy, and recommended quality cuts to maximize sample purity and completeness. This catalog is supplemental to the DLA NN and GP Finder VAC.
@@
ZLyA VAC contains the redshifts and broad absorption line (BAL) information used for the DR1 Ly forest BAO analysis (DESI Collaboration et al., 2025a) The catalog was constructed with main-survey, dark-program quasars with a Redrock ZWARN flag of either zero or four (see §3.1.2).525252https://desidatamodel.readthedocs.io/en/latest/bitmasks.html#zwarn, Redshifts were updated from the LSS catalog redshifts to correct for a systematic underestimation of quasar redshifts introduced by improper handling of Lyman-series optical depth (Brodzeller et al., 2023; Bault et al., 2025). The new redshifts were determined using a modified version of the Redrock HIZ quasar templates that incorporate the Kamble et al. (2020) model for Ly effective optical depth redshift evolution. BAL attributes were derived using the method presented in Filbert et al. (2024).

### v2

@@ F.2 Milky Way Survey VACs
MWS VAC provides analysis of stellar spectra by the Milky Way Survey Working Group (koposov25a). The analysis involves fitting DESI spectra using two pipelines: RVSpecFit (koposov24a) and FERRE (allende-prieto23a) (see Cooper et al. 2023 and koposov24a for more details). The RVSpecFit pipeline derives radial velocity measurements, stellar parameters, and abundances, while FERRE provides stellar parameters and abundance measurements. Both pipelines have been applied to approximately 6.5M coadded spectra of stars or possible stars from DESI DR1. Additionally, the RVSpecFit pipeline was run on approximately 9.5M single-exposure spectra, enabling radial velocity variability analyses. Since the DESI EDR, both pipelines have been significantly improved, with RVSpecFit now using neural networks to interpolate stellar spectra, thus eliminating issues with clustering of stellar parameters on grid nodes that was present in the EDR VAC. The data products in this VAC include tables of stellar parameters, radial velocities, their uncertainties, and the best-fit models for the DESI spectra. In addition, the data tables have been cross-matched with the latest Gaia DR3 release.
@@
MWS BHB VAC contains spectroscopically confirmed BHB stars. BHB stars are excellent tracers of the Milky Way halo; they are intrinsically bright, so they can be observed to large distances, and there exists a simple polynomial relationship between their color and absolute magnitude (deason11a; belokurov16a), meaning that their distances can be computed with high precision. BHB candidates were targeted in the MWS (see §4.4.4 in Cooper et al., 2023) and as secondary targets observed in dark time (see Myers et al., 2023). Due to photometric uncertainties, targeted BHB candidates contain contamination primarily from more nearby blue straggler stars. The DESI spectroscopic observations, however, allow us to remove this contamination. The BHB VAC contains over 6,300 spectroscopically confirmed BHB stars, all having distances derived from photometry, as described in bystrom25a. This BHB catalogue contains the stellar parameters from the RVSpecFit pipeline (koposov24a), including radial velocity, surface gravity and effective temperature, and reaches heliocentric distances of about 120 kpc.
@@
MWS SpecDis VAC provides distance measurements for over 4M stars from DR1 following the selection criteria outlined in li25a. The VAC is based on a data-driven approach which adopts a neural network to establish the connection between stellar spectra and luminosity. It is trained on Gaia parallax and Gaia -band apparent magnitude measurements, with the following key improvements: (1) We adopt a special training label to avoid cutting off negative parallaxes; (2) We do not include any parallax uncertainty cuts for the training sample, but we include the parallax uncertainties in the loss function. This approach ensures that the training sample does not suffer from biases in parallax, and our training sample includes a significant number of distant giants. We can achieve significant improvements in our measurements than Gaia parallaxes beyond 7 kpc. (3) We adopt principal component analysis to decrease the noise and the dimensionality of the spectra. In general, the distance uncertainties decrease with increasing average signal-to-noise ratio in the B and R arms of the DESI stellar spectra (see Table 2). We adopt a Gaussian mixture model to identify candidate binary stars. The final catalog offers distance, distance uncertainty, the binary candidate flag plus other useful photometric and astrometric information, cross-matched from the RVSpecFit values reported in the MWS VAC (koposov24a) and Gaia DR3.
@@
SPDist VAC provides spectrophotometric distances in the form of absolute magnitude in the Gaia -band for all stars observed by the DESI MWS, using a data-driven approach similar to thomas22a. Distances are estimated using a fully connected multi-layer perceptron (MLP) model using Gaia DR3 photometry (gaia-collaboration23a), spectroscopic metallicity from the SP pipeline, and atmospheric parameters (effective temperature and surface gravity) derived from a combination of values from both the RV and SP pipelines (koposov24a) as inputs. The MLP outputs the absolute magnitude in the Gaia -band, which is subsequently used to calculate distances. Some stars exhibit significant discrepancies in spectroscopic parameters derived by the RV and SP pipelines, likely leading to unreliable distance estimates. These suspicious stars are flagged with the FLAG_GOOD parameter. We recommend using this flag along with the quality flags recommended for the RV and SP pipelines (koposov24a). For distance modulus computations, we recommend using the median of the absolute magnitude distribution (MG_50), while uncertainties can be derived from the 16th and 84th percentiles of the distribution. These uncertainties should be combined in quadrature with the intrinsic precision of the method (0.167 mag, 8% relative distance precision) to obtain the total uncertainty. The primary difference between the methodologies of SpecDist and SPdist is that SpecDist predicts distances from the stellar spectra, whereas SPdist predicts distances from a full list of stellar parameters by the MWS RV and SP pipelines. A forthcoming paper will present a detailed comparison of the two methods.
@@
Stellar Reddening VAC contains the observed spectra and the RVSpecFit model spectra of stars that provided reddening measurements for the DESI dust map (zhou24a). The DESI dust map, publicly released with zhou24a, is used in the DESI DR1 cosmology analysis (DESI Collaboration et al., 2024d). Also included in this VAC is the stellar catalog containing the per-star reddening measurements in the Dark Energy Camera (DECam; flaugher15a) filters as well as other properties from the DESI pipeline and the imaging catalogs. While the paper uses the first two years of DESI data, only the DR1 portion of the data is released in the VAC. Also note that while both the MWS VAC and this VAC use RVSpecFit and are based on the same DESI data, there are some key differences: (1) the RVSpecFit model spectra in this VAC are the zero-extinction spectra, whereas the MWS VAC model spectra include an additional multiplicative term to fit the smooth-varying component of the observed spectra; and (2) the DR1 MWS VAC uses a different (newer) version of RVSpecFit than the version used here. For general purposes (other than for studying dust), the MWS VAC should be used instead of this VAC.
@@ F.3 Extragalactic Science VACs
HETDEX VAC contains HETDEX (hill21a; gebhardt21a) and DESI spectra of HETDEX-selected Ly emitter candidates (mentuch-cooper23a) followed up by DESI. The VAC also contains the emission line fits from both sets of spectra. Details of the analysis can be found in landriau25a.
@@
DESIVAST VAC contains cosmic voids identified within the DESI DR1 volume. The void locations are computed using a volume-limited subsample of the BGS Bright sample (see Hahn et al. 2023) extending to . We apply evolutionary corrections to 479,486 galaxies and enforce a magnitude cut of . We then apply three void-finding algorithms: VoidFinder (el-ad97a; hoyle02a; douglass22a), V2/VIDE (neyrinck08a; sutter15a), and V2/REVOLVER (nadathur19a), to obtain three comparable void catalogs. Our VoidFinder catalog contains 3,765 voids, with 1,489 non-edge voids defined as those not bordering on the survey edges. For V2/VIDE (V2/REVOLVER), we find 1,478 (1,992) voids with 297 (389) non-edge voids. Further information regarding this VAC can be found in rincon25a.
@@
Dwarf Galaxy VAC presents a sample of extragalactic dwarf galaxies () identified in DESI DR1. The catalog includes galaxies selected from the BGS and ELG samples, as well as the LOW-Z secondary target program (darragh-ford23a), spanning a redshift range of . However, due to differences in target selection criteria across these samples, the catalog is not uniformly selected. Stellar masses are derived using CIGALE525252https://cigale.lam.fr (boquien19a; siudek24a) and optical color-based prescriptions (mao24a; de-los-reyes24a). In a forthcoming paper (Manwadkar et al. 2026, in preparation), we will supplement this catalog with information on spectroscopic and photometric completeness, and an analysis of quenched fractions as a function of environment.
@@
EmFit VAC provides emission-line fitting results using the EmFit code (pucha25a) run on low-redshift () galaxy spectra, focusing on eight nebular emission lines and doublets: H, [O III] 4959,5007, [N II] 6548,83, H, and [S II] 6716,31. The code tests for the presence of extra components in the [S II] and [O III] doublets independently, and the profiles of the rest of the narrow components are designed to match the profile of the [S II] lines. It further tests for the presence of a possible broad component in the Balmer lines. The flux and width measurements along with their uncertainties for all detected components are reported in this catalog (pucha25a).
@@
FastSpecfit VAC delivers a broad range of observed-frame, rest-frame, and intrinsic physical parameters for all extragalactic () targets in DR1 observed by DESI (moustakas23b; Moustakas et al. 2026, in preparation).535353https://fastspecfit.readthedocs.io It uses physically motivated stellar population synthesis and emission-line templates to model each DESI spectrum jointly with the optical through infrared broadband photometry. The measurements included in this VAC include: stellar velocity dispersions; stellar masses; K-corrections and rest-frame colors and magnitudes; emission-line velocity widths, velocity shifts, fluxes, and equivalent widths (separately both narrow and broad emission-line components); and much more.
@@
Extended Halo-based Group VAC is the updated version of the extended halo-based group catalog derived from the Legacy Surveys DR9 for a galaxy sample limited to -band apparent magnitude . For all the DESI target classes included in DR1, the photometric redshifts will be updated to spectroscopic redshifts. yang21a provide more details regarding the extended halo-based group finder and sample selection.
@@
Mass EMLines VAC provides stellar mass and emission line measurements for all galaxies in DESI DR1 with reliable redshift measurements. Stellar masses are derived using CIGALE (boquien19a), which employs the broadband -, -, -, -, and -band photometry from the Legacy Surveys, and spectrophotometry of 10 artificial bands generated through convolution with DESI spectra. A main set of optical emission lines are measured by a single Gaussian fit, with absorption correction through stellar continuum fitting performed using STARLIGHT545454http://www.starlight.ufsc.br (cid-fernandes05a; cid-fernandes11a). Additionally, the catalog includes stellar population properties derived by CIGALE and those obtained by STARLIGHT using DESI spectra. For additional information see zou24b and Zou et al. (2026, in preparation).
@@
Strong Lensing VAC provides a comprehensive catalog of spectroscopic observation of strong gravitational lenses observed by DESI. The observations are primarily obtained through a program that seeks to obtain spectroscopic redshifts for both the lensing galaxies and the lensed sources for 1800 new strong lensing candidate systems (huang25a; Storfer et al. 2025, in preparation). Objects included in the catalog were visually inspected in order to make a determination on the quality of the Redrock redshift, and to determine the nature of the system. Systems with at least one confirmed redshift for either the lens or source are included in the VAC. In addition to relevant data from the default DESI pipeline, the VAC includes data from the FastSpecFit VAC, including velocity dispersion, star-formation rate, and stellar mass. Also included is diameter aperture photometry at the location of the fibers targeting lenses and sources. Lastly, with a number of Ly emitters and absorbers not currently fit for redshifts via Redrock, we use custom template-fitting within the Redrock framework to fit for our own redshifts. Systems that are assigned redshifts this way are noted in the catalog.
@@ F.4 Quasar Science VACs
AGN/Galaxy Classification VAC provides AGN and QSO identification for galaxies () from all target classes observed within DR1 for objects with at least one DESI spectrum. Identification of QSOs and AGN is achieved through multiple means, including optical and ultraviolet emission-line diagnostics and multi-wavelength photometry. The VAC enables rapid selection of more complete versus more pure methodologies, as well as selection through any one AGN diagnostic (Juneau et al. 2026, in preparation). A tutorial accompanies the catalog which demonstrates both how to use the catalog for rapid AGN / QSO selection as well as using the open source diagnostic code to select a specific line diagnostic and to choose the dividing line methodologies of those diagnostics. An example might be to test the [S II] baldwin81a classification assuming the more recent dividing lines of law21a compared with the traditional dividing lines of kewley01b, which were reported to potentially overestimate the presence of AGN (e.g., pucha25a).
@@
AGN Host Properties VAC provides physical properties for approximately 17M galaxies derived through spectral energy distribution fitting using CIGALE (v.22.1; boquien19a). This analysis accounts for contributions from both stars and, if present, an AGN. CIGALE uses the principle of energy balance, where dust-absorbed stellar emission in the ultraviolet and optical bands is re-emitted in the infrared; it estimates key galaxy and AGN properties such as stellar mass, star formation rate, and the AGN fraction (the contribution of the AGN dusty torus to the total IR luminosity (e.g., ciesla15a; salim18a; yang20a; yang22a). A detailed description of the catalog and its statistical properties based on data in the DESI EDR can be found in siudek24a, with additional follow-up analysis using spectral energy distribution modeling to identify AGN to be described in siudek25a.
@@
BHMass VAC provides Mg II-based supermassive black hole mass estimates for 490,648 quasars at from DESI DR1. Approximately 35% of the sample have mass uncertainties less than 0.5 dex. The iron-corrected mass is estimated using the fitting process and estimator introduced by pan25a, which accounts for the effects of the Eddington ratio. Additionally, we integrate fitting results from the FastSpecFit VAC and offer alternative mass estimates based on the formulations by shen11a; le20a; yu23a, providing users with flexibility in choosing their preferred estimator. This catalog will be periodically updated and extended to include more DESI quasars.
C IV Absorber VAC contains the largest catalog to date of C IV absorption systems detected in quasar spectra from DESI DR1 (anand25a). The catalog includes C IV absorber systems identified along quasar sightlines over the redshift range . Absorbers are detected using an automated matched-kernel convolution technique combined with adaptive signal-to-noise thresholds. Quasar continua are modeled using a non-negative matrix factorization (NMF) approach. Rest-frame equivalent widths are measured for both doublet components using a double-Gaussian profile fitting method, and column densities are derived using the apparent optical depth method (savage91a). Catalog completeness is quantified via Monte Carlo absorber-injection simulations, with the sample reaching approximately 50% completeness at , while the overall catalog purity exceeds 95%.
@@
DLA NN and GP Finder VAC contains the catalog of DLAs used for the DR1 Ly BAO analysis. The catalog results from the combination of two DLA finding algorithms. The first one is based on a convolution neural network trained with mock DESI spectra whose algorithm and performance is described in wang22a. The second algorithm, which uses a Gaussian Process (GP), is based on the GP model of ho20a. Both outputs from these algorithms are retained in the combined catalog and DLA candidates are merged when they are found by both methods with a velocity difference of less than 800 km s-1. The merged catalog includes estimated redshifts and H I column densities together with confidence flags from both algorithms. It contains 54,416 candidate DLAs with H I column density cm-2 within Ly forest spectra with signal-to-noise ratio larger than three.
@@
DLA Template Finder VAC contains candidate DLAs identified with the DLA Toolkit software, which uses spectral template-fitting. The catalog was constructed from quasar spectra with , allowing for a maximum of three detections per sight-line; it contains an estimated DLA redshift, HI column density, and detection significance for each DLA. brodzeller25a provide the full details regarding the construction of the catalog, parameter accuracy, and recommended quality cuts to maximize sample purity and completeness. This catalog is supplemental to the DLA NN and GP Finder VAC.
@@
ZLyA VAC contains the redshifts and broad absorption line (BAL) information used for the DR1 Ly forest BAO analysis (DESI Collaboration et al., 2025b) The catalog was constructed with main-survey, dark-program quasars with a Redrock ZWARN flag of either zero or four (see §III.1.2).555555https://desidatamodel.readthedocs.io/en/latest/bitmasks.html#zwarn, Redshifts were updated from the LSS catalog redshifts to correct for a systematic underestimation of quasar redshifts introduced by improper handling of Lyman-series optical depth (Brodzeller et al., 2023; Bault et al., 2025). The new redshifts were determined using a modified version of the Redrock HIZ quasar templates that incorporate the kamble20a model for Ly effective optical depth redshift evolution. BAL attributes were derived using the method presented in filbert24a.